I'm having trouble with parallel and perpendicular lines that contribute with slope. Here's what it is asking from me:
"Write an equation in slope-intercept form of the line that passes through the given point and is parallel to the graph on the given equation."
The given equation is (1,-3); y+2= 4*(x-1)
Thank you for your help in advance!

Answer by richwmiller(17219)

y=mx+b is the slope intercept form
(1,-3); y+2= 4*(x-1)
y+2= 4*(x-1) is almost in that form
y=4x-4-2
y=4x-6
slope is 4
a line parallel will have the same slope of 4
-3=4*1+b
-3=4+b
-7=b
y=4x-7 is line that is parallel and passes through (1,-3)